The Flower City Grocery is faced with the following budgeting decision. Its display freezer system should be repaired. But to repair may not be the correct approach. Rather the company could purchase a new freezer system for $15,000. The new freezer system has more display space and will increase profits. With the new system net cash inflows will average $8,000 and cash operating costs will be $3,000 annually. The company has a cost of capital of 9 percent. The tax rate for the company, thanks to the recent Republican ead tax bill is only 25%. The new freezer system will last 5 years, with no maintenance expected. The company depreciates assets using the straight-line method over the asset's useful life.

Required:

1. Compute the net present value of this new freezer system. Is the system acceptable?

2. What is the internal rate of return of this new freezer system? Is the system acceptable?

3. What is the payback period for this new freezer system? If the company has a required payback period of three years, is this system acceptable?

4. If straight-line depreciation is used for this system, what is the accounting simple rate of return?
Answer #1
1]
a] Initial investment $           15,000
Net cash inflows $             8,000
-Cash operating costs $             3,000
-Depreciation [15000/5] $             3,000
=Net operating income $             2,000
-Tax at 25% $                500
=NOPAT $             1,500
+Depreciation $             3,000
=Annual operating cash flow $             4,500
b] PV of annual OCF = 4500*(1.09^5-1)/(0.09*1.09^5) = $           17,503
Less: Initial investment $           15,000
NPV $             2,503
As the NPV is positive, the system is acceptable.
2] IRR is that discount rate for which the NPV is 0, which
means that PV of cash inflows = Initial investment.
Hence,
15000 = 4500*PVIFA(r,5), where r is the
IRR.
Solving for r
PVIFA(r,5) = 15000/4500 = 3.3333
From the annuity factor tables, the factors
for 15% and 16% are 3.3522 and 3.2743
respectively.
Hence, IRR falls between 15% and 16%.
By simple interpolation, IRR = 15%+1%*(3.3522-3.3333)/(3.3522-3.2743) = 15.24%
AS the IRR is greater than the cost of capital of 9%, the
system is acceptable.
3] Payback period = 15000/4500 = 3.33 Years
As the payback of the system is more than the maximum
acceptable of 3 years, the system is not acceptable.
4] Accounting simple rate of return = 1500/(15000/2) = 20.00%